A taxidriver filled his car petrol tank with 40 litres of petrol on Monday. The next day, he filled the tank with 50 litres of petrol. If the petrol costs ₹ 44 per litre, how much did he spend in all on petrol?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Petrol quantity filled on Monday = 40 litres
Petrol quantity filled on Tuesday = 50 litres
Total petrol quantity filled = (40 + 50) litre
Cost of petrol per litre =₹44
Total money spent = 44 × (40 + 50)
= 44 × 90
=₹ 3960

Thanks